A force of 2.30×104 NN is applied to a diver's face mask that has an area of 115 cm2cm2 .

Part A

Find the pressure in atmatm on the face mask. (1 atmatm = 101,325 PaPa)

Express your answer with the appropriate units.
Answer #1

we have

F = 2.30 * 10^4 N

and

A = 115 cem^2 = 115 * (0.01)^2 = 0.0115 m^2

we know,

P = F / A

or

P = 2.30 * 10^4 / 0.0115 = 2.0 * 10^6 N / m^2 = 2.0 * 10^6 Pa = 2000 KPa = 2000 / 101.325 = 19.7 atm

thus

pressure = 19.7 atm